If (x-2) and (x+2) are factors of the expression x^3+ax^2+bx+1 what is the sum of a and b
1
Expert's answer
2021-08-27T09:57:52-0400

Given "x-2" is a factor of the expression "x^3+ax^2+bx+1"


"x-2=0"

"x=2"

Substituting 2 in the equation "x^3+ax^2+bx+1"


"2^3+a(2^2)+b(2)+1=0"

"8+4a+2b+1=0"

"4a+2b=-9........................i"


Given "x+2" is a factor of the expression "x^3+ax^2+bx+1"


"x+2=0"

"x=-2"

Substituting -2 in the equation "x^3+ax^2+bx+1"


"-2^3+a(-2^2)+b(-2)+1=0"

"-8+4a-2b+1=0"

"4a-2b=7........................ii"

Solving equation i and ii by substitution

"4a+2b=-9........................i"

"4a-2b=7.........................ii"



"4a=7+2b"

Hence "7+2b+2b=-9"

"7+4b=-9"

"4b=-16"

"b=-4"


"4a+2(-4)=-9"

"4a=-1"

"a=-\frac{1}{4}"


The sum of a and b will be

"a+b = -4-\frac{1}{4}=-\frac{17}{4}"